For questions 1 and 2, please remember that there are two replication forks per origin, and that the synthesis of the two new daughter strands occurs at the same rate and at the same time.

1. (4 points) The pathogenic strain O157:H7 of E. coli has one circular chromosome. The DNA molecule in this chromosome contains 5,498,450 bp. DNA synthesis at one replication fork occurs at a rate of 70,000 nucleotides per minute.

How many minutes will it take to completely replicate the molecule? (Round to the nearest whole number): Please show your calculations in the space below.

2. (5 points) A bacterial cell synthesizes DNA at one replication fork at a rate of 700 nucleotides per second. If the cell replicates its DNA molecule in 20 minutes, how many base pairs are found in its DNA molecule?

Please show your calculations in the space below.

For questions 3 and 4, please remember that only one RNA strand is being synthesized, and only the template strand of the DNA is being transcribed.

3. (3 points) E. coli RNA polymerase adds ribonucleotides to an RNA strand at a rate of 22 per second. The size of the transcribed region of the largest gene in E. coli is 1538 bp. How long (in seconds) will it take RNA polymerase to complete transcription of an average-sized gene? (Round off to the nearest whole number):

Please show your calculations in the space below.

4. (4 points) The average size of the transcribed region in a human gene is 10,000 bp. If a eukaryotic RNA polymerase adds ribonucleotides at a rate of 28 per sec, how long (in minutes) will it take the RNA polymerase to complete transcription of an average human gene? (Round off to the nearest whole number):

Please show your calculations in the space below.
Answer #1

Ans 1. Size of the chromosome = 5,498,450 bp.

Replication fork nucleotides addition rate is 70,000 bp/ min.

But there are two replication fork in DNA replication so, the total nucleotides added by them is 140,000 bp/min.

Total time taken for chromosome synthesis = chromosome size / rate of synthesis by replication fork.

= 5,498,450 / 140,000 = 39.27 min = 39 min.

Thus, 39 min is required to synthesize the chromosome.

Ans 2:- Replication fork movement rate is 700 my / sec.

Now, there are two replication synthesis nt in 1 min = 2 * 700 * 60 = 84,000 bp / min.

So, the total genome size is = 20 min * 84,000 bp /min. = 1,680,000 bp.

Thus, the bacterial cell have 1,680,000 bp DNA.

Ans 3:- RNA polymerization rate = 22 nt / sec.

The average gene size is 1538 bp.

Thus, the time taken to synthesize is 1538/ 22 = 69.9 sec or 70 sec.

So, 70 sec RNA polymerase require to synthesize average gene.

Ans 4 :- Average size of transcribed gene is 10,000 bp.

RNA polymerase rate is 28 nt /sec or 1680 bp/min.

Total time taken to synthesize gene = 10,000 / 1680 = 5.95 min or 6 min.

Thus, RNA polymerase will take 6 min to synthesise the average gene.

Dear student if you find the answer convincing please do give the feedback.